# One Language, Many Gaps: Evaluating Dialect Fairness and Robustness of Large Language Models in Reasoning Tasks

### v2

@@ 4.2 AAVE Conversational Norms
We use the mutual information between the token distributions of SE and AAVE ReDial to find that the top 5 most informative AAVE features in terms of distinguishing them from SE are , up, in, gon, and gotta. We note that many of these features are not well-known AAVE-specific features (e.g., up). Through a further investigation of our dataset, we find that these lexicons are associated with phrase-level AAVE constructions. For instance, instead of saying …encode the answer… in Standard English, AAVE instruction says …wrap it up…. This finding is particularly interesting as we find that, in addition to previous linguistic observations of AAVE-specific morphosyntactic features, there are conversational norms of the dialect such as nuanced uses of phrases.
@@
We compute Spearman’s correlation between the frequency of the features we find in each instance and their corresponding performance drop. Indeed, these features play a significant role in predicting GPT-4o’s performance degradation (r=-0.318, p<0.001). We further implement and analyze 12 rule-based AAVE features following Ziems et al. (2022) (details in Appendix A.11), which are well documented in linguistic literature such as "finna" as a maker of immediate future Nguyen and Grieve (2020). We notice that the influential lexical features are a subset of the feature set discovered by mutual information (i.e., some of the actual influential features are not encoded in synthetic transformation rules). Consequently, the influence of synthetic features is not as strong as those discovered by mutual information (r=-0.256, p<0.001). This means that simple rule-based transformations that implement the most salient morphosyntactic AAVE features may not be able to capture rich, context-dependent use of the dialect and, therefore fall short in predicting LLMs’ performance in real workflows.
@@
Aided by GPT-o1 preview to filter the vast amount of data, we conducted a linguistically informed analysis of LLMs’ most frequent errors on AAVE. For the algorithmic tasks, grammatical constructions and non-standard verb forms (e.g., finna’, ’em), omission of articles and auxiliary verbs may cause the model to misinterpret references and function naming conventions. For example, GPT-4o interprets you gon’ write a python function, python_function as a general statement rather than a directive to name the function. On logic tasks, the frequent use of double negatives, zero copula, and inverted conditionals introduces structural ambiguities. For example, the construct He don’t take no breaks can invert intended meanings, leading the model to misunderstand conditional statements. On math, informal expressions, unclear quantity references, and non-standard comparative cause erroneous parsing of numerical information and confusion over collective versus individual quantities. Informal phrasings like half as much as he be runnin’ and ambiguous comparative expressions (4 fewer boxes of apple pie than on Sunday) can cause the model to misinterpret numerical relationships, resulting in errors when calculating totals, differences, or fractions. On the integrated task, phonetic spellings, colloquial connectors, and inverted word orders limit the model’s understanding of concurrency, and stepwise instructions. Such dialectal nuances highlight the necessity of our dataset and also call for more efforts to collect more human data for relevant purposes.
@@ 5 Related Work
Previous works on AAVE studies in natural language processing mostly focus on non-reasoning-heavy tasks such as POS tagging (Jørgensen et al., 2015, 2016), language identification and dependency parsing (Blodgett et al., 2016), automatic captioning (Tatman, 2017), and general language generation (Deas et al., 2023). AAVE is also found to be more likely to trigger false positives in hate speech identifiers (Davidson et al., 2019; Sap et al., 2019) due to specific word choices (Harris et al., 2022), be considered negative by automatic sentiment classifier (Groenwold et al., 2020), and cause covert biases in essential areas of social justice (Hofmann et al., 2024). Relevant studies (Ziems et al., 2022; Gupta et al., 2024) also find that rule-based AAVE feature perturbations can downgrade language model performance in a wide range of tasks covered by GLUE (Wang, 2018).
More generally, dialects in world languages pose challenges to natural language processing systems. Ziems et al. (2023) find that auto-encoder models are brittle on rule-based English dialect feature perturbations. Fleisig et al. (2024) report that English dialect speakers perceive responses generated by chatbots to be more negative than Standardized English (SE) prompts. Faisal et al. (2024) find that world dialects cause problems in tasks including dependency parsing (Scherrer et al., 2019) and machine translation (Mirzakhalov, 2021) on mBERT and XLM-R (Conneau et al., 2020).
Fairness and Robustness of Large Language Models. LLMs are widely testified to be both unfair and brittle. They introduce unfair performance (Huang et al., 2023; Dong et al., 2024) and cost (Petrov et al., 2024) to users across different languages, exacerbate social imbalance by marginalizing minority groups in various aspects including gender (Kotek et al., 2023; Fraser and Kiritchenko, 2024), race (Hofmann et al., 2024; Wang et al., 2024), and culture (Naous et al., 2023; Tao et al., 2024). They also show different performance to reasoning tasks in different languages (Huang et al. 2023, 2024; Ranaldi et al. 2024; inter alia). For the first time, our study provides extensive empirical evidence that LLMs exhibit unfairness in reasoning tasks. This bias specifically affects speakers of certain dialects within a single language.
Previous works report that LLMs are very sensitive to slight variations of prompts by introducing typos or paraphrasing in SE (Elazar et al., 2021; Liang et al., 2022; Raj et al., 2022; Zhu et al., 2023b; Lin et al., 2024). In this work, we consider a novel application of using human-written perturbations in AAVE by asking humans to rewrite instances to their dialect and evaluate LLM robustness towards these natural perturbations, which have proven to cause LLMs to be more brittle than synthetic typo-style (Section 4.1) or linguistic-rule-based (Appendix A.9) perturbations.
@@ 6 Conclusion
Our study is the first attempt to objectively evaluate the dialect robustness and fairness of LLMs across reasoning tasks. We introduce ReDial, a dataset comprising over 1.2K parallel prompts in Standardized English and African American Vernacular English (AAVE) tailored to reasoning tasks: algorithm, logic, math, and integrated reasoning. Extensive empirical evidence on ReDial demonstrates that LLMs exhibit significant unfairness and brittleness when reasoning tasks are expressed in AAVE. These findings underscore the unfairness to dialect users and LLMs’ brittleness with natural prompt variations with the same semantics. We advocate for further research to enhance dialect fairness and robustness of LLMs, ensuring equal service for all linguistic groups and demographics.
@@ 7 Limitations
First, as the first systematic framework for analyzing LLM bias in dialectal queries for reasoning tasks, we selected AAVE due to its linguistic significance and cultural impact. However, we recognize the vast diversity of dialects worldwide. The insights derived from AAVE may not generalize to other dialects. To ensure annotation quality and maintain the focus of our study, we concentrated on AAVE with high-quality human annotations. Future research could expand on our framework to encompass a wider range of dialects, generating more broadly applicable conclusions.
Second, our benchmark, ReDial, evaluates LLM performance across four categories of reasoning tasks using queries sampled from seven popular and well-documented benchmarks. While these tasks are representative of common reasoning challenges, we acknowledge that reasoning is a multifaceted domain with many additional categories and tasks that fall outside the scope of this study.
Third, we evaluated five representative LLM families in this study, including widely used and sota models. However, given the rapid proliferation of new LLMs, testing every model is infeasible. We hope that future research will use the ReDial benchmark to investigate fairness and reasoning robustness across a broader range of LLMs as they emerge.
@@ 8 Ethic Statement
We do not collect annotators’ personal information; while we firmly commit to this rule to protect annotators’ privacy, it makes it difficult to draw conclusions about how annotators’ backgrounds shape their writing/individual-level variations. Further on the ethical aspect of data collection, we work with a data vendor that makes sure the recruitment and annotation adhere to high standards for and from the annotators. However, although we have a legal contract and we try our best to convey our guidelines and requirements, we admit that we do not have full control over how the vendor recruits people and conducts data annotation.
@@ 9 Acknowledgement
We thank all the bodies who have provided funding for the authors and for the associated project. FL is supported by Clarendon and Jason Hu studentship. ELM is supported by the Alan Turing Institute. We thank Ambrosio Blanco for helpful feedback on the ethical review. We are grateful to the people who offered feedback and suggestions along the way, and in particular, Janet B. Pierrehumbert, Su Lin Blodgett, Wenshan Wu, Xiaoyuan Yi, Yan Xia and other colleagues in Microsoft Research, who offered invaluable advice and helped us refine the paper.
@@ A.3.1 Employment Information
We work with data vendors to employ 13 annotators in total for our task. For algorithm instance annotation, we specifically hire annotators with computer science backgrounds. Annotators are self-identified as proficient speakers of African American Vernacular English. We do not pose any hard constraints in verifying dialect identity as previous studies do (e.g., Ziems et al. (2023)). We note even within a dialect there can be significant variations on the individual level and that we want to avoid homogenization and over-simplification of the dialect (King, 2020). Instead, we ask self-identified annotators to cross-check each other’s annotations and modify if they sound unnatural.
@@
Information Collected We do not collect personally identifiable information from our annotators (e.g., name, age, etc). We only collect the annotators’ responses to our consent form and their annotations of our data.
@@
No AI Assistant We explicitly inform our annotators that they should not reply on any AI assistant tools to help them complete the task. To further ensure this, we design our annotation platform to disallow copy and paste. The default annotation area for annotators is the original text, which means that it is easier for annotators to simply edit the text than querying AI assistants.
@@ A.4 Data Quality Verification
You will be given two prompts, one in Standard English and one in African American English. Determine whether the African American English prompt is a valid paraphrase of the Standard English prompt. Ignore the semantic validaty of the Standard English prompt. Standard English: "[SAE_PROMPT]" African American English: "[AAVE_PROMPT]" Is the African American English prompt a valid paraphrase of the Standard English prompt?
@@ A.6 Results for Non-zero Temperature
A.7 Full Results on ReDial
@@
A.8 Perplexity of Typos vs. Natural AAVE Text
We gradually inject typos in SE ReDial by replacing/deleting/adding words/characters, such that we make the input texts more difficult for LLMs (i.e., the measured perplexity goes up). We find that while these perturbations degrade model performance, the drop did not reach the severity observed with natural AAVE data on large-scale models (see full results in Section A.8). This discrepancy suggests that AAVE brittleness is not solely due to the general difficulty in processing. Interestingly, larger models such as LLaMA-3.1-70B-Instruct and Phi-3-Medium-128K-Instruct degrade less on synthetic noise yet still struggle disproportionately with authentic AAVE data. This suggests their brittleness stems from reliance on SE priors and limited encoding of AAVE-specific discourse structures.
A.9 Multivalue Perturbation
A.10 Qualitative Analysis of Standardized Prompting
We observe a sensible gap between model performance on Standardized ReDial with zero-shot prompting and AAVE ReDial with standardization prompting. In this section, we qualitatively compare GPT-4o’s outputs in these two settings, the model with the best absolute overall performance, and examine its errors. We focus on the math subset of ReDial and identify three key error patterns: wrong question rephrasing, distraction by irrelevant information, and failure to execute all steps.
Wrong question rephrasing. GPT-4o wrongly phrases question ‘Jame … How many years have they got between them now if in 8 years his cousin will be 5 years younger than twice his age?’ to ‘James … How old is his cousin now?’, which changes the question of age gap to absolute age.
Distraction by irrelevant information. GPT-4o gets distracted by task-irrelevant information after AAVE standardization while the distraction is not observed in Standardized ReDial. For instance, in ‘Say we got 8 different books and 10 different movies in the crazy silly school series. How many more movies than books is there gon be in the crazy silly school series if you read 19 books and watched 61 movies?’, books that have been read and movies that have been watched are not associated with the answer. Although GPT-4o can ignore irrelevant information in Standardized ReDial, it gets distracted after AAVE standardization, which shows the brittleness of its reasoning ability.
Failure to execute all the steps. GPT-4o sometimes simulates an algorithm to solve math problems after standardization (e.g., ‘Let $(x)$ be the number of apple pie boxes…’). However, it does not fully solve the problem in the end and only returns a formula (e.g., ‘30x + 255’), which indicates that the model’s reasoning ability is limited when it comes to program simulation for queries expressed in dialects.
A.11 Synthetic Lexical Feature list
A.12 Statement of Contribution

### v3

@@ 4.2 AAVE Conversational Norms
We use the mutual information between the token distributions of SE and AAVE ReDial to find that the top five most informative AAVE features in terms of distinguishing them from SE are , up, in, gon, and gotta. Note that many features here are not well-known AAVE-specific features (e.g., up). Through a further investigation of our dataset, we find that these lexicons are associated with phrase-level AAVE constructions. For instance, instead of saying …encode the answer… in SE, AAVE instruction says …wrap it up…. This finding is particularly interesting because, in addition to previous linguistic observations of AAVE morphosyntactic features Sidnell (2002); Martin and Wolfram (2013), there are important conversational norms of the dialect, such as nuanced uses of phrases Green (2002); Morgan (2002); Rickford and Rickford (2007).
@@
We compute Spearman’s correlation between the frequency of the features we find with mutual information in each instance and their corresponding performance drop. Indeed, these features play a significant role in predicting GPT-4o’s performance degradation (r=-0.318, p<0.001). We further implement and analyze 12 rule-based AAVE features following Ziems et al. (2022) (details in Appendix A.9), which are well documented in linguistic literature such as finna as a marker of immediate future Nguyen and Grieve (2020). We notice that the influential lexical features are a subset of the feature set discovered by mutual information (i.e., some of the actual influential features are not encoded in synthetic transformation rules). Consequently, the influence of synthetic features is not as strong as those discovered by mutual information (r=-0.256, p<0.001). This means that simple rule-based transformations that implement the most salient morphosyntactic AAVE features may not be able to capture rich, context-dependent use of the dialect and, therefore, fall short in predicting LLMs’ performance in real workflows.
@@
With assistance from GPT-o1 preview to filter the vast amount of data, we conducted a linguistic analysis of frequent errors in AAVE. For algorithm tasks, grammatical constructions and non-standard verb forms (e.g., finna’, ’em), omission of articles and auxiliary verbs may cause the model to misinterpret references and function naming conventions. For example, GPT-4o interprets you gon’ write a python function, python_function as a general statement rather than a directive to name the function. On logic tasks, the frequent use of double negatives, zero copula, and inverted conditionals introduces structural ambiguities (e.g., He don’t take no breaks). On math, informal expressions, unclear quantity references, and non-standard comparative constructions cause erroneous parsing of numerical information and confusion over collective versus individual quantities. Informal phrasings like half as much as he be runnin’ and ambiguous comparative expressions (4 fewer boxes of apple pie than on Sunday) can cause the model to misinterpret numerical relationships. On the integrated task, phonetic spellings, colloquial connectors, and inverted word orders limit the model’s ability to understand concurrency and follow stepwise instructions. Such dialectal nuances highlight the necessity of our dataset and also call for more efforts to collect more human data for relevant purposes. [^7] We also provide more analysis comparing reasoning chains in AAVE/SE ReDial in Section A.8.
@@ 5 Related Work
Previous work on AAVE studies in natural language processing mostly focuses on non-reasoning-heavy tasks such as POS tagging (Jørgensen et al., 2015, 2016), language identification and dependency parsing (Blodgett et al., 2016), automatic captioning (Tatman, 2017), and language modeling (Deas et al., 2023). AAVE is found to be more likely to trigger false positives in hate speech identifiers (Davidson et al., 2019; Sap et al., 2019) due to word choices (Harris et al., 2022), to be considered negative by automatic sentiment classifier (Groenwold et al., 2020), and to cause covert biases in essential areas of social justice (Hofmann et al., 2024). Other studies (Ziems et al., 2022; Gupta et al., 2024) find that rule-based AAVE perturbations can downgrade language model performance in GLUE (Wang, 2018).
More generally, dialects across world languages pose challenges to natural language processing systems. Ziems et al. (2023) find that auto-encoder models are brittle on rule-based English dialect feature perturbations. Fleisig et al. (2024) report that responses generated by chatbots to dialectal inputs are perceived as more negative by English dialect speakers than responses to SE prompts. Faisal et al. (2024) find that non-standardized dialects cause problems in dependency parsing (Scherrer et al., 2019) and machine translation (Mirzakhalov, 2021) on mBERT and XLM-R (Conneau et al., 2020).
Fairness and Robustness of Large Language Models. LLMs exhibit both unfairness and brittleness. They offer unfair performance (Huang et al., 2023; Dong et al., 2024) and cost burdens (Petrov et al., 2024) to users of different languages, marginalize minority groups across dimensions such as gender (Kotek et al., 2023; Fraser and Kiritchenko, 2024), race (Hofmann et al., 2024; Wang et al., 2024; Sun et al., 2025), and culture (Naous et al., 2023; Tao et al., 2024). They also show different performance to reasoning tasks in different languages (Huang et al. 2023, 2024; Ranaldi et al. 2024; inter alia). To our knowledge, our study provides the first extensive empirical evidence that LLMs are unfair in reasoning tasks. This bias specifically affects speakers of certain dialects within a single language.
Previous works report that LLMs are brittle when prompts are varied through typos or paraphrasing in SE (Elazar et al., 2021; Liang et al., 2022; Raj et al., 2022; Zhu et al., 2023b; Lin et al., 2024; Röttger et al., 2024). We use a novel approach of human-written perturbations in AAVE and evaluate LLM robustness towards these natural perturbations, which result in greater brittleness than synthetic typo-style (Section 4.1) or linguistic-rule-based (Appendix A.7) perturbations.
@@ 6 Conclusion
Our study is the first to objectively evaluate the dialect robustness and fairness of LLMs in reasoning. We introduce ReDial, a dataset of over 1.2K parallel prompts in Standardized English and African American Vernacular English (AAVE) tailored to algorithm, logic, math, and integrated reasoning. Extensive empirical evidence on ReDial demonstrates that LLMs exhibit significant unfairness and brittleness when reasoning tasks are expressed in AAVE. These findings underscore the unfairness to dialect users and LLMs’ brittleness with natural prompt variations with the same semantics. We advocate for further research to enhance dialect fairness and robustness of LLMs, ensuring equal service for all linguistic groups and demographics.
@@ 7 Limitations
First, as the first systematic framework for analyzing LLM bias in dialectal queries for reasoning tasks, we selected AAVE due to its linguistic significance and cultural impact. However, we recognize the vast diversity of dialects worldwide. The insights derived from AAVE may not generalize to other dialects. To ensure annotation quality and maintain the focus of our study, we concentrated on AAVE with high-quality human annotations. Also, we only have one annotator to annotate each instance and another AAVE annotator for naturalness check due to the limited budget. While we try to ensure diversity by hiring more annotators (13 vs. 3 in previous literature; Ziems et al., 2022) and spreading annotation over more instances, we are aware that this might still bring subjectivity into our benchmark. Future research could expand on our framework to encompass a wider range of dialects, hiring a more diverse range of annotators, and generating more broadly applicable conclusions.
Second, our benchmark, ReDial, evaluates LLM performance across four categories of reasoning tasks using queries sampled from seven popular and well-documented benchmarks. While these tasks are representative of common reasoning challenges in both fundamental (e.g., the logic tasks) and practical areas (e.g., code generation and complex planning for multi-agent systems in the integrated tasks), we acknowledge that reasoning is a multifaceted domain with many additional categories and tasks that fall outside the scope of this study (e.g., medical/financial reasoning).
Third, we evaluated five representative LLM families in this study, including widely used and state-of-the-art models. However, given the rapid proliferation of new LLMs, testing every model is infeasible. We hope that future research will use the ReDial benchmark to investigate fairness and reasoning robustness across a broader range of LLMs as they emerge.
Fourth, due to the difficulty of gathering large-scale dialect data for training, we cannot perform additional analysis on how supervised fine-tuning/reinforcement learning with relevant data might help models bridge the dialect gap, which makes it difficult to draw conclusions about how different training methods affect dialect robustness. Although it might be technically possible to generate high-quality synthetic data (with known difficulties discussed in Sections 1 and 4), we consider it to be out of the scope of our paper. Despite so, we do observe that general supervised fine-tuning/reinforcement learning do not bridge the gap as can be observed in comparing models in the GPT family. We hope that future research can develop a reliable and scalable way to gather more high-quality dialect data.
@@ 8 Ethic Statement
We do not force annotators to disclose their personal information; while we firmly commit to this rule to protect annotators’ privacy, it makes it difficult to draw detailed conclusions about how annotators’ backgrounds shape their writing/individual-level variations. Further on the ethical aspect of data collection, we work with a data vendor that makes sure the recruitment and annotation adhere to high standards for and from the annotators. However, although we have a legal contract and we try our best to convey our guidelines and requirements, we admit that we do not have full control over how the vendor recruits people and conducts data annotation.
@@ 9 Acknowledgement
We thank all the bodies who have provided funding for the authors and for the associated project, including the OpenAI Research Access credits to FL. FL is supported by Clarendon and Jason Hu studentship. ELM is supported by the Alan Turing Institute. We thank Ambrosio Blanco for helpful feedback on the ethical review. We are grateful to the people who offered feedback and suggestions along the way, and in particular, Su Lin Blodgett, Wenshan Wu, Xiaoyuan Yi, Yan Xia, Jing Yao, Sunayana Sitaram, and other colleagues in Microsoft Research, who offered invaluable advice and helped us refine the paper.
@@ A.3.1 Employment Information
We work with data vendors to employ 13 annotators in total for our task. For algorithm instance annotation, we specifically hire annotators with computer science backgrounds. Annotators are self-identified as proficient speakers of African American Vernacular English. We do not pose any hard constraints in verifying dialect identity as previous studies do (e.g., Ziems et al. (2023)). We note even within a dialect there can be significant variations on the individual level, and that we want to avoid homogenization and over-simplification of the dialect (King, 2020). Instead, we ask self-identified annotators to cross-check each other’s annotations and modify if they sound unnatural.
We only have one annotator for each prompt due to the budget limitation. However, we ensure diversity and reduce subjectivity in two ways for each annotation. First, we ensure diversity by hiring more annotators to annotate more instances (as opposed to one annotator per instance, which may introduce biases). We report the average score across datasets; this allows us to provide a diverse range of reference points, as the overall diversity is preserved by spreading annotations across many instances. For example, having 1,600 instances annotated by 4 annotators (with each annotator covering 400 instances) achieves a similar level of diversity as having 400 instances where all 4 annotators annotate every instance. After annotating for each instance, another AAVE speaker, who was not the annotator for that instance, ensures the resulting translation is consistent and sounds natural.
@@
Information Collected We do not force disclosure of personal information from our annotators (e.g., name, age, etc). We only make it mandatory that we collect the annotators’ responses to our consent form and their annotations of our data.
Demographic information We report information on those (11 annotators) willing to disclose more demographic information. Annotators’ ages range between 23 and 35 years old, with 2 female and 9 male annotators. 4 of them have Master’s degrees, and others have Bachelor’s degrees.
@@
No AI Assistant We explicitly inform our annotators that they should not reply on any AI assistant tools to help them complete the task. To further ensure this, we design our annotation platform to disallow copy and paste. The default annotation area for annotators is the original text, which means that it is easier for annotators to simply edit the text than to query AI assistants.
@@ A.4 Data Quality Verification
You will be given two prompts, one in Standard English and one in African American English. Determine whether the African American English prompt is a valid paraphrase of the Standard English prompt. Ignore the semantic validaty of the Standard English prompt. Standard English: "[SE_PROMPT]" African American English: "[AAVE_PROMPT]" Is the African American English prompt a valid paraphrase of the Standard English prompt?
@@ A.6 Results for Non-zero Temperature
A.7 Multivalue Perturbation
A.8 Qualitative Analysis of Reasoning Chains
We qualitatively compared GPT-4o’s reasoning chain in SE and AAVE ReDial. We focus on the math subset of ReDial and identify two key error patterns: (1) distraction by irrelevant information and (2) failure to execute all steps. Below, we briefly sketch our findings.
Distraction by irrelevant information. GPT-4o gets distracted by task-irrelevant information in AAVE ReDial, while we do not observe the same in SE ReDial. For instance, in ‘Say we got 8 different books and 10 different movies in the crazy silly school series. How many more movies than books is there gon be in the crazy silly school series if you read 19 books and watched 61 movies?’, books that have been read and movies that have been watched are not associated with the answer. Although GPT-4o ignores irrelevant information in SE ReDial prompts, it cannot do so in AAVE, showing its reasoning ability’s brittleness.
Failure to compose the program and execute all steps. GPT-4o sometimes simulates an algorithm to solve math problems. However, it gets stuck in one step of its reasoning chain where it confuses the reference of comparison. For instance, in a question ’On Saturday, he sold … 4 fewer boxes of apple pie, than on Sunday. Come Sunday he done sold 5 more boxes of gingerbread than on Saturday and 15 more boxes of apple pie.’ GPT-4o reasons by starting with ’Let $A_{s}$ be the number of boxes of apple pie sold on Saturday…Let $A_{u}$ be the number of boxes of apple pie sold on Sunday’. Then, it realizes there is a contradiction between the correct formula $A_{s}=A_{u}-4$ and the wrong formula $A_{u}=A_{s}+15$ , where it wrongly considers the number of apple pies sold on Sunday to be 15 more than that on Saturday. This indicates that reasoning with queries expressed in dialects limits a model’s reasoning ability.
A.9 Synthetic Lexical Feature list
A.10 Statement of Contribution
A.11 Full Results on ReDial